20120927

LUCI Flex Calib not yet done.

Glycol leak prevented today's tests

LUCI GEIRS would not run on obs2-5 workstations. Font conflict in X11 corrected by SHooper after phone call

LUCI AGw not initializing. After some tests from Tucson,
  • Power cycled by yanking the plug up on AGw1
  • Power cycle did not work, AGw1 down for now.

DGH says we can open. Fairly uniform cirrus across sky. Switching to MODS for pointing and collimation.0

UT 05:45 Enclosure is open, MODS is awake.
R0001, B0001 sieveSnap

### DETXY dice5 repeats to check AGw transform UT 06:09 execMODS StoneO_1136_act0_5dicexy5.img 30s exposures
  • R and B images are 0002-0026

### Re-measure rotator center for pointing model UT 7:07 Re-measure rotator center, execMODS each once:
UT=== R- B- Script Name
07:07 27 27 StoneO_1136_trk0.img
07:11 28 28 StoneO_1136_trk180.img
07:16 29 29 StoneO_1136_trk270.img
07:20 30 30 StoneO_1136_trk90.img

### RADEC dice5 repeats to check AGw transform UT 07:25 execMODS StoneO_1136_act0_5dice5.img 30s exposures
  • R and B images are 0031-0038
  • Script aborted...clouds.

UT 07:53 Dome closed

### Closed-dome LUCI calibrations
  • luci1.20120927.0001 10s o2dcr w/lamp off for "dark"
  • turn on AR lamp
  • rotate to -90 to start
  • 0002 is a test exposure

FC_N180_150Ks.cal @EL=90 0003 - 0034 (20 and 55 degrees done by MEdwards this afternoon)

##### Re-opening UT 09:14

### Take pointing model data UT 10:01
  • Turn on Pointing LOG!
  • Use /home/LBTO/POINTING/Pointingstars5, watch for fields with EL<25!
  • Use AllStars10Faint (MODS catalog) w/star on axis to preset
Filter on AGw is: clear

repeat {
  • preset
  • collimate
  • rangebal
  • note subsequent WFS #s
  • Match guide # to WFS using TIME-OBS
}

Bin  StarID      WFS
  -  ACT_0xxF    279  
  1  WT10_368F   289
  2  ACT_0409F   - 25d and rising, revisit
  3  WT10_405F   294
  4  ACT_0340F   300
  5  WT10_274F   - 22d, skip
  6  WT10_189F   309 - G* lost in clouds at end
  7  WT10_263F   315
  8  WT10_076F   -   too low
  9  WT10_178F   322
 10  WT10_267F   329
 11  ACT_0218F   -   too low
 12  WT10_333F   335
 13  -           -   no matched star
 14  ACT_0303F   341
 15  WT10_252F   345
 16  WT10_299F   351
 17  ACT_0195F   357
 18  WT10_093F   -   too low
 19  ACT_0140F   362
 20  WT10_314F   370
 20' ACT_0282F   375
 21  WT10_244F   381
 22  WT10_206F   -   too low
 23  WT10_293F   390
  2  ACT_0409F   393
 24  ACT_0329F   400
 25  ACT_0395F   406
 26  WT10_408F   -   too low
 27  WT10_407F   410
 28  ACT_0344F   415
 29  WT10_381F   422
 30  WT10_374F   428
 33  WT10_368F   436
<verbatim>

Turn off Pointing LOG!  File in /home/telescope/PointingLogs/20120927_?X.log

WFS images to use for collimation model: ~LBTO/wfs_collim_data.tar.gz <br>
List of wfs files to use for collimation model: ~LBTO/wfs_20120927.lis <br>

I looked at the data quickly and it mostly looks good.  

-- Main.DavidThompson - 27 Sep 2012</verbatim>
<nop>
Topic revision: r1 - 27 Sep 2012, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback